### v2

# Almost Global Asymptotic Trajectory Tracking for Fully-Actuated Mechanical Systems on Homogeneous Riemannian Manifolds

Abstract
In this work, we address the design of tracking controllers that drive a mechanical system’s state asymptotically towards a reference trajectory. Motivated by aerospace and robotics applications, we consider fully-actuated systems evolving on the broad class of homogeneous spaces (encompassing all vector spaces, Lie groups, and spheres of any finite dimension). In this setting, the transitive action of a Lie group on the configuration manifold enables an intrinsic description of the tracking error as an element of the state space, even in the absence of a group structure on the configuration manifold itself (e.g., for $\mathbb{S}^{2}$). Such an error state facilitates the design of a generalized control policy depending smoothly on state and time, which drives the geometric tracking error to a designated origin from almost every initial condition, thereby guaranteeing almost global convergence to the reference trajectory. Moreover, the proposed controller simplifies elegantly when specialized to a Lie group or the $n$-sphere. In summary, we propose a unified, intrinsic controller guaranteeing almost global asymptotic trajectory tracking for fully-actuated mechanical systems evolving on a broad class of manifolds. We apply the method to an axisymmetric satellite and an omnidirectional aerial robot.
I Introduction
An efficient, effective approach to the control of mechanical systems is to synthesize an overall controller via the composition of a low-rate “open-loop” planner (to select a reference trajectory) with a high-rate “closed-loop” tracking controller (to drive the system towards the reference trajectory). Such layered control architectures are found across an extremely diverse array of engineered and biological systems, demonstrating the effectiveness of the paradigm [2].
For some mechanical systems, it is possible to reduce the “tracking problem” to the easier “regulation problem” (i.e., asymptotically stabilizing an equilibrium) for the same system, using a state-valued “tracking error” between the reference and actual states that “vanishes” only when the two are equal (e.g., $({q-q_{d}},{\dot{q}-\dot{q}_{d}})$ for a mechanical system on ${\mathbb{R}^{n}}$). Then, feedback is designed to drive the error towards the origin, i.e., a constant (equilibrium) trajectory. Such a reduction is substantially easier for fully-actuated (vs. underactuated) systems, since arbitrary forces can be exerted, both to compensate for time variation in the reference and to inject suitable artificial potential and dissipation terms to almost globally asymptotically stabilize the error [3].
Clearly, a smooth tracking controller suitable for operation on the entire state space (in general, the tangent bundle of a non-Euclidean manifold) cannot rely on a local, coordinate-based tracking error. When the configuration manifold is a Lie group, [4] shows that the group structure furnishes an intrinsic, globally-defined tracking error on the state space, via the action of the group on itself. For this reason, [4] remarks that “the tracking problem on a Lie group is more closely related to tracking on $\mathbb{R}^{n}$ than it is to the general Riemannian case, for which the group operation is lacking.”
Tracking via error regulation on manifolds that may not be Lie groups is proposed in [1]. The authors present a theorem ensuring almost global tracking for fully-actuated systems on compact Riemannian manifolds, assuming the control inputs can be chosen to solve a “rank one” overdetermined linear system at each state and time. Solutions exist in some special cases (e.g., on $\mathbb{S}^{2}$ with carefully chosen potential, dissipation, and kinetic energy), although they are non-unique. In other cases, no solutions exist, and the question of when solutions exist in general is left open. Also, the inputs may not depend continuously on state and time, a desirable property that justifies seeking only almost global asymptotic stability [5].
Figure 1: The proposed controller, applied to a mechanical system on $\mathbb{S}^{2}$ (e.g., the axisymmetric satellite). We show snapshots of parallel rollouts from a random sampling of initial states (configuration and velocity) in $T\mathbb{S}^{2}$. All sampled rollouts converge to the reference trajectory.
In view of these observations, and in pursuit of a systematic approach to tracking control for fully-actuated mechanical systems in yet a broader setting, a question emerges:
On which class of manifolds can the tracking problem be reduced to the regulation problem?
We pose this question in a global or almost global sense; if purely local convergence suffices, the answer is trivial (i.e., “all of them”), since smooth manifolds are locally Euclidean.
In this work, we consider mechanical systems evolving on homogeneous spaces, a class of manifolds that includes all Lie groups but is more general. We obtain globally valid “error dynamics” by extending an intrinsic error used in tracking for kinematic systems in [6] to the mechanical setting, exploiting the transitive action of a Lie group on the configuration manifold even when it lacks a group structure of its own. Ultimately, we contribute a systematic approach to tracking control with formal guarantees for a broad class of systems that includes many space, aerial, and underwater robots.
II Mathematical Background
For any map ${(x,y)\mapsto f(x,y)}$, we define the map ${f_{x}:y\mapsto f(x,y)}$ (resp. ${f^{y}:x\mapsto f(x,y)}$) by holding the first (resp. second) argument constant. For a smooth map ${f:M\to N}$ between manifolds, its derivative is denoted ${\operatorname{d}f:TM\to TN}$, while the dual of its derivative is the unique map ${\operatorname{d}f^{*}:T^{*}N\to T^{*}M}$ such that ${\big{\langle}{\hskip 1.0pt}\omega_{n}{\hskip 1.5pt;\hskip 1.5pt}\operatorname{d}f(v_{m}){\hskip 1.0pt}\big{\rangle}=\big{\langle}{\hskip 1.0pt}\operatorname{d}f^{*}(\omega_{n}){\hskip 1.5pt;\hskip 1.5pt}v_{m}{\hskip 1.0pt}\big{\rangle}}$ for any vector ${v_{m}\in TM}$ and any covector ${\omega_{n}\in T_{f(m)}^{*}N}$. The derivative of a function ${V:M\to\mathbb{R}}$ induces a covector field ${\operatorname{d}V:M\to T^{*}M}$. We assume all given curves are smooth.
II-A Homogeneous Riemannian Manifolds
Definition 1 (see [7, Ch. 3]).
A homogeneous Riemannian manifold $(Q,\Phi,\kappa)$ is a smooth manifold $Q$ equipped with:
  1. i.
    a transitive left action $\Phi:G\times Q\to Q$ of a Lie group $G$, i.e., ${\forall\ q_{1},q_{2}\in Q}$, ${\exists\ g\in G\textrm{ s.t. }\Phi(g,q_{1})=q_{2}}$, and
  2. ii.
    a $\Phi$-invariant Riemannian metric $\kappa$, i.e., ${\forall\ g\in G}$ and ${v_{q},w_{q}\in TQ},\ \kappa\big{(}{v_{q}},{w_{q}}\big{)}=\kappa\big{(}{\operatorname{d}\Phi_{g}(v_{q}),\operatorname{d}\Phi_{g}(\omega_{q})}\big{)}$.
The stabilizer $G_{q}$ of ${q\in Q}$ is the maximal subgroup $G$ such that ${\forall\ g\in G_{q},\ \Phi(g,q)=q}$. The identity element is ${\mathds{1}\in G}$.
Example 1 (name=The $n$-Sphere,label=example:sphere).
For any ${n\in\mathbb{N}}$, consider the sphere $\mathbb{S}^{n}$ and the rotation group ${SO(n+1)}$. For notational convenience, we model these manifolds as embedded submanifolds of Euclidean spaces, i.e., we let ${\mathbb{S}^{n}=\{s\in\mathbb{R}^{n+1}:s^{\textrm{T}}s=1\}}$ and we let ${SO(n+1)=}$${\{R\in\mathbb{R}^{n+1\times n+1}:R^{\textrm{T}}R=R\,R^{\textrm{T}}=\mathds{1},\ \det R=1\}}$. Let $\rho$ and $\Psi$ be the respective restrictions to ${\mathbb{S}^{n}\subset\mathbb{R}^{n+1}}$ of the Euclidean metric and of the usual action of ${SO(n+1)}$, i.e.,
$\begin{gathered}{\Psi}:SO(n+1)\times\mathbb{S}^{n}\to\mathbb{S}^{n},\ (R,q)\mapsto R\,q.\end{gathered}$ (1)
Then, it can be shown that $(\mathbb{S}^{n},\Psi,\rho)$ is an ${SO(n+1)}$-homogeneous Riemannian manifold. Moreover, $\rho$ induces the canonical isomorphisms ${\rho^{\flat}:T\mathbb{S}^{n}\to T^{*}\mathbb{S}^{n},v_{q}\to v_{q}{}^{\textrm{T}}\textrm{ and}}$ ${\rho^{\sharp}:T^{*}\mathbb{S}^{n}\to T\mathbb{S}^{n},f_{q}\to f_{q}{}^{\textrm{T}}}$, and ${G_{q}\cong SO(n)}\ \forall\ {q\in\mathbb{S}^{n}}$.
Example 2 (name=A Lie Group,label=example:group).
Consider any Lie group $G$, the left action of $G$ on itself (i.e., ${L:(h,g)\mapsto hg}$), and any $L$-invariant metric $\kappa_{\mathbb{I}}$. Then by [8, Thm. 5.38], there exists an inner product $\mathbb{I}$ on ${\mathfrak{g}\cong T_{\mathds{1}}G}$ such that ${\forall\ v_{g},w_{g}\in TG}$,
$\kappa_{\mathbb{I}}(v_{g},w_{g})=\mathbb{I}\,\big{(}\operatorname{d}L_{g^{-1}}(v_{g}),\operatorname{d}L_{g^{-1}}(w_{g})\big{)}.$ (2)
Moreover, $(G,L,\kappa_{\mathbb{I}})$ is a homogeneous Riemannian manifold. $L$ is a free action, i.e., for ${g\neq\mathds{1}}$, the map $L_{g}$ has no fixed points. Thus, for any ${g\in G}$, the stabilizer is just ${G_{g}=\{\mathds{1}\}}$. It is clear that for all ${\xi\in\mathfrak{g}}$ and ${\tau\in\mathfrak{g}^{*}}$, we have ${\kappa_{\mathbb{I}}^{\flat}(\xi)=\mathbb{I}^{\flat}(\xi)}$ and ${\kappa_{\mathbb{I}}^{\sharp}(\tau)=\mathbb{I}^{\sharp}(\tau)}$ (see, e.g., [8, Sec. 2.3.4]).
II-B Navigation Functions
The following functions are described in [3], in the more general setting of manifolds with boundary. They are useful in generating artificial potential forces for stabilization.
Definition 2.
On a boundaryless manifold $Q$, a $0_{Q}$-navigation function is a proper Morse function ${P:Q\to\mathbb{R}}$ that has the point ${0_{Q}\in Q}$ as its unique local minimizer. A navigation function is said to be perfect if its domain admits no other navigation function with fewer critical points.
For example, for any ${0_{\mathbb{S}^{n}}\in\mathbb{S}^{n}}$ and $k_{P}>0$, the map
$P_{\mathbb{S}^{n}}:\mathbb{S}^{n}\to\mathbb{R},\ q\mapsto-k_{P}\,({0_{\mathbb{S}^{n}}})^{\textrm{T}}\,q$ (3)
[…]
$P_{SE(3)}:\begin{pmatrix}R&x\\ 0_{1\times 3}&1\end{pmatrix}\mapsto\operatorname{tr}\big{(}K_{R}(\mathds{1}-R)\big{)}+x^{\mathrm{T}}K_{x}\,x,$ (4)
where ${x\in\mathbb{R}^{3}}$, ${R\in SO(3)}$, and $K_{x},K_{R}$ are $3\times 3$ symmetric positive-definite matrices where $K_{R}$ has distinct eigenvalues. Navigation functions on $\mathbb{R}^{3}$, $SE(2)$, $SO(3)$, and ${SO}(2)$ can be obtained from (4), and a navigation function on the product of boundaryless manifolds can be given by the sum of navigation functions on each factor [9]. Thus, we can easily obtain an explicit navigation function on any product space whose factors are all $n$-spheres or closed subgroups of $SE(3)$, capturing most homogeneous spaces of interest.
II-C Fully-Actuated Mechanical Systems
Definition 3.
A fully-actuated mechanical system on a homogeneous Riemannian manifold ${(Q,\Phi,\kappa)}$ has dynamics
$\nabla_{\dot{q}}\dot{q}=\kappa^{\sharp}(f_{q}),$ (5)
[…]
Remark 1 (Generality).
Even if the system were subject to additional state-dependent forces ${F:TQ\to T^{*}Q}$, or it were governed by the Riemannian connection $\widetilde{\nabla}$ of a different metric $\widetilde{\kappa}$ that fails to be $\Phi$-invariant, the system can be rendered in the form (5) by the static state feedback
$f_{q}(\dot{q})$ $=\widetilde{\kappa}^{\flat}\big{(}\big{(}{\widetilde{\nabla}-{\nabla}}\big{)}({\dot{q}},{\dot{q}})-F(\dot{q})+\tilde{\kappa}^{\sharp}({f}_{q}^{\hskip 0.5pt\prime})\big{)},$ (6)
where $\big{(}{\widetilde{\nabla}-{\nabla}}\big{)}$ is the “difference tensor” (see [7, Prop. 4.13]) between the Riemannian connection of $\tilde{\kappa}$ and of an invariant metric ${\kappa}$, while ${f}_{q}^{\hskip 0.5pt\prime}$ is a “virtual” input. Since any homogeneous space with compact stabilizer admits an invariant metric [7, Cor. 3.18], a tracking controller for a $\Phi$-invariant unforced system will, in practice, yield a tracking controller for the more general system via (6). Also, the invariant case is common, and it simplifies the computations.
Definition 4.
A control policy ${(t,\dot{q})\mapsto f_{q}(t,\dot{q})}$ for the system (5) achieves almost global asymptotic tracking of a reference trajectory ${q_{d}:\mathbb{R}\to Q}$ if, for each ${t_{0}\in\mathbb{R}}$, there exists a residual set ${S_{t_{0}}\subseteq TQ}$ of full measure such that if ${\dot{q}(t_{0})\in S_{t_{0}}}$, then ${\operatorname{dist}_{TQ}\big{(}\dot{q}(t),\dot{q}_{d}(t)\big{)}\to 0}$ as ${t\to\infty}$. Note that $\operatorname{dist}_{TQ}$ is the Riemannian distance in $TQ$ corresponding to the Sasaki metric (the natural Riemannian metric on $TQ$ induced by $\kappa$).
III An Intrinsic, State-Valued Tracking Error
In this section, we take an approach similar to that of [6] (which dealt with kinematic systems) to describe an intrinsic state-valued tracking error suitable for mechanical systems.
III-A Tracking Error in Homogeneous Spaces
Definition 5.
For an origin ${0_{Q}}$ in $(Q,\Phi,\kappa)$ and a smooth curve ${q_{d}:\mathbb{R}\to Q}$, a $0_{Q}$-lift of $q_{d}$ is any smooth curve ${g_{d}:\mathbb{R}\to G}$ such that $\Phi\big{(}g_{d}(t),0_{Q}\big{)}=q_{d}(t)\textrm{ for all }t\in\mathbb{R}.$
Definition 6.
For any actual and reference configuration trajectories ${q,q_{d}:\mathbb{R}\to Q}$ and a given $0_{Q}$-lift $g_{d}$ of $q_{d}$, the configuration error trajectory is the smooth curve given by
$e:\mathbb{R}\to Q,\ t\mapsto\Phi\big{(}{g_{d}}(t)^{-1},q(t)\big{)},$ (7)
while the state error trajectory is its derivative, ${\dot{e}:\mathbb{R}\to TQ}$.
In [6], the authors use a tracking error of the form (7) for kinematic (i.e., first-order) systems to synthesize optimal tracking controllers with local convergence. (In fact, such an error state has its roots in observer design [10].) We lift this definition to the tangent bundle to obtain a state-valued tracking error for mechanical (i.e., second-order) systems. The following extends [6, Prop 4.1] to the second-order setting.
Proposition 1.
Consider any actual and reference trajectories ${q,q_{d}:\mathbb{R}\to Q}$. Let ${g_{d}:\mathbb{R}\to G}$ be a $0_{Q}$-lift of ${q_{d}}$, and let ${0_{TQ}\in TQ}$ be the zero tangent vector at $0_{Q}$. Then, for any time ${t\in\mathbb{R}}$, ${\dot{q}(t)=\dot{q}_{d}(t)}$ if and only if ${\dot{e}(t)=0_{TQ}}$.
Proof.
The actual and reference configuration can be expressed (suppressing the $t$-dependence for brevity) as
$q$ $=\Phi(g_{d},\Phi({g_{d}^{-1}},q))=\Phi({g_{d}},e),\quad q_{d}=\Phi(g_{d},0_{Q}),$ (8)
since by assumption, $g_{d}$ is a $0_{Q}$-lift of $q_{d}$. Thus, we have
$\dot{q}=\operatorname{d}\Phi^{e}(\dot{g}_{d})$ $+\operatorname{d}\Phi_{g_{d}}(\dot{e}),$ (9a)
$\dot{q}_{d}=\operatorname{d}\Phi^{0_{Q}}(\dot{g}_{d})$ $+\operatorname{d}\Phi_{g_{d}}(0_{TQ}).$ (9b)
Assuming for sufficiency that ${\dot{e}=0_{TQ}}$ (and therefore ${e=0_{Q}}$), (9a)-(9b) immediately implies that ${\dot{q}=\dot{q}_{d}}$. Assuming for necessity that ${\dot{q}=\dot{q}_{d}}$ (and therefore ${q=q_{d}}$), we recall that for each ${g\in G}$, the maps $\Phi_{g}$ and $\operatorname{d}\Phi_{g}$ are automorphisms of $Q$ and $TQ$ respectively. For this reason, (8) implies ${e=0_{Q}}$, while that conclusion and (9a)-(9b) imply that ${\dot{e}=0_{TQ}}$. ∎
In summary, the intrinsic tracking error smoothly transforms the state space (in a manner depending smoothly on time) in such a way that only the current reference state is mapped to the zero tangent vector over the origin of the lift. Also, for the ($\Phi$-invariant) Riemannian distance ${\operatorname{dist}_{Q}}$, (8) implies that ${\operatorname{dist}_{Q}\big{(}q(t),q_{d}(t)\big{)}=\operatorname{dist}_{Q}\big{(}e(t),0_{Q}\big{)}}$. Thus, (7) encodes the distance from the reference configuration “accurately”.
III-B Computing Horizontal Lifts of Curves
The following notions will aid in computing a lift of a given reference trajectory (and ultimately, the tracking error).
Definition 7 (see [11, Sec. 23.4]).
[…]
In fact, every homogeneous Riemannian manifold admits a (perhaps non-unique) reductive decomposition [12, Prop. 1].
Proposition 2 (Horizontal Lifts in Reductive Homogeneous Spaces).
Consider a homogeneous Riemannian manifold $(Q,\Phi,\kappa)$ with origin ${0_{Q}\in Q}$ and reductive decomposition ${\mathfrak{g}=\mathfrak{f}\oplus\mathfrak{q}}$. For each smooth curve ${q_{d}:\mathbb{R}\to Q}$ and initial point ${g_{0}\in G}$ such that ${\Phi(g_{0},0_{Q})=q_{d}(0)}$, there exists a unique smooth curve ${g_{d}:\mathbb{R}\to G}$ (called the horizontal lift of $q_{d}$ through $g_{0}$) such that (i) ${g_{d}}$ is a $0_{Q}$-lift of $q_{d}$, (ii) $g_{d}(0)=g_{0}$, and (iii) $\operatorname{d}L_{g_{d}(t)}^{-1}\big{(}\dot{g}_{d}(t)\big{)}\in\mathfrak{q}$ for all $t\in\mathbb{R}$.
Proof.
By [13, Prop 9.33], $G$ is a (left) principal $F$-bundle over ${Q=G/F}$. In particular, the projection map is given by ${\pi:G\to Q,\ g\mapsto\Phi(g,0_{Q})}$, while the free and proper action is given by ${\Upsilon:F\times G\to G,(f,g)\mapsto R_{f^{-1}}(g)}$. Using the $\operatorname{Ad}_{F}$-invariance of $\mathfrak{q}$, it can be shown that ${HG=}$${\{\operatorname{d}L_{g}(\mathfrak{q}):g\in G\}}$ is a principal connection on ${\pi:G\to Q}$. Then, the claim follows from the existence and uniqueness of horizontal lifts in principal bundles (see [14, Sec. 2.9]). ∎
Remark 2 (Sections, Lifts, and Nontrivial Bundles).
The proof of Proposition 2 describes a sense in which a lift projects “down” to the original curve via $\pi$. If the principal bundle ${\pi:G\to Q}$ is trivial (i.e., ${G\cong Q\times F}$ globally and not merely locally), then there exist global sections ${\sigma:Q\to G}$, i.e., smooth maps satisfying ${\pi\circ\sigma=\operatorname{id}}$. In fact, any such section furnishes a (perhaps non-horizontal) lift ${g_{d}:t\mapsto\sigma\circ q_{d}(t)}$. However, when ${\pi:G\to Q}$ is a nontrivial bundle (e.g., the bundle corresponding to $(\mathbb{S}^{2},\Psi,\rho)$, namely ${\pi:SO(3)\to\mathbb{S}^{2}}$), the nonexistence of global sections makes it impossible to generate global lifts using a section. Nor can the initial value $g_{0}$ of a horizontal lift depend continuously on $q_{d}(0)$ alone. Thus, continuous deformation of the reference trajectory can result in discontinuous changes in the tracking error. However, such a discontinuity is with respect to the choice of reference trajectory (i.e., the planning layer); once a reference trajectory $q_{d}$ has been selected, the configuration error will depend smoothly on both time and state. Note also that even for a horizontal lift $g_{d}$ of some $q_{d}$, when ${q_{d}(t_{1})=q_{d}(t_{2})}$ it may still be that ${g_{d}(t_{1})\neq g_{d}(t_{2})}$ due to nontrivial “holonomy” (see [14, Fig. 3.14.2]), which cannot always be avoided.
Remark 3 (Computing Horizontal Lifts Numerically).
Let ${(\pi\times\varphi^{i}):}$ ${\pi^{-1}(U_{i})\subseteq G}$$\to$} ${(U_{i}\subseteq Q)\times F,}$ ${i\in\{1,\ldots,k\}}$ be a collection of local trivializations covering ${\pi:G\to Q}$. Let ${\mathbb{A}^{i}:TU_{i}\to\mathfrak{f}}$ be the local connection forms for $HG$ (see [14, Prop. 2.9.12]). Suppose ${q_{d}(t)\in U_{j}}$ for ${t\in[t_{1},t_{2}]}$ and let ${f^{j}:[t_{1},t_{2}]\to F}$ solve the initial value problem (IVP)
$f^{j}(t_{1})=\varphi^{\hskip 0.75ptj}\big{(}g(t_{1})\big{)},\quad\dot{f}^{j}(t)=\operatorname{d}L_{f^{j}(t)}\circ\mathbb{A}^{j}\big{(}\dot{q}_{d}(t)\big{)}.$ (10)
Then the horizontal lift $g_{d}$ through $g_{0}$ of the curve $q_{d}$ satisfies
$g_{d}(t)=(\pi\times\varphi^{\hskip 0.75ptj})^{-1}\big{(}q_{d}(t),f^{j}(t)\big{)}\textrm{ for all }t\in[t_{1},t_{2}].$ (11)
Moreover, the restriction of the reference trajectory to any finite interval may be subdivided into finitely many segments, each contained within a single trivialization. We repeatedly solve the IVP (10) numerically for each segment, ultimately reconstructing a smooth lift in $G$ via (11). Computationally, this is preferable to solving an IVP in $G$ directly, since it ensures that integration error will accumulate only along the fibers (vs. horizontally). Hence, numerical integration accuracy is not paramount (reducing the computational burden), since the solution will be an exact lift, even if it is not perfectly horizontal. The value of the lifted reference at each time $t$ can also be computed “just in time” to compute the tracking error.
Example 3 (name=The $n$-Sphere,continues=example:sphere).
On the homogeneous Riemannian manifold $(\mathbb{S}^{n},\Psi,\rho)$, we choose the origin ${0_{\mathbb{S}^{n}}=(0,\ldots,1):=e_{n}}$ and identify $\mathfrak{so}(n)$ with the ${n\times n}$ skew-symmetric matrices. Following [11, Sec. 23.5], we have the reductive decomposition ${\mathfrak{so}(n+1)=\mathfrak{f}\oplus\mathfrak{q}}$, where
$\mathfrak{f}=\left\{\begin{pmatrix}\xi&0\\ 0&0\end{pmatrix}:\xi\in\mathfrak{so}(n)\right\},\ \mathfrak{q}=\left\{\begin{pmatrix}0&\upsilon\\ -\upsilon^{\textrm{T}}&0\end{pmatrix}:\upsilon\in\mathbb{R}^{{n}}\right\}.$
We horizontally lift a given configuration reference trajectory ${q_{d}:\mathbb{R}\to\mathbb{S}^{n}}$ in the sense of Proposition 2 to obtain the lifted reference ${R_{d}:\mathbb{R}\to{SO(n+1)}}$, performing numerical computations in the manner of Remark 3 for accuracy. Then, the configuration error in the form (7) is simply given by
$e:\mathbb{R}\to\mathbb{S}^{n},\ t\mapsto{R_{d}}(t)^{\textrm{T}}\,q(t).$ (12)
Example 4 (name=A Lie Group,continues=example:group).
Since the stabilizer of any point $g$ on the homogeneous Riemannian manifold $(G,L,\kappa_{\mathbb{I}})$ is ${G_{g}=\{\mathds{1}\}}$, $0_{G}$-lifts (of any kind) are unique for each ${0_{G}\in G}$. Making the usual choice ${0_{G}=\mathds{1}}$, the lift is the original curve itself, and the configuration error is
$e:\mathbb{R}\to G,\ t\mapsto{g_{d}}(t)^{-1}\,g(t),$ (13)
showing that in the special case of Lie groups, the configuration error reduces elegantly to a familiar, intuitive form, called the “right group error function” [8, p. 548]. In the additive group $G=(\mathbb{R}^{n},+)$ (i.e., a vector space), the configuration error is (unsurprisingly) the map ${t\mapsto g(t)-g_{d}(t)}$.
IV Almost Global Asymptotic Tracking
We now present the main result on trajectory tracking. First, we define the body velocity of any curve ${\gamma:\mathbb{R}\to G}$ to be the curve in $\mathfrak{g}$ given by ${t\mapsto\operatorname{d}L_{\gamma(t)}^{-1}\big{(}\hskip 0.5pt\dot{\gamma}(t)\hskip 0.5pt\big{)}}$.
Theorem 1.
Consider a fully-actuated mechanical system on $(Q,\Phi,\kappa)$, a reference trajectory ${q_{d}:\mathbb{R}\to Q}$, and any ${0_{Q}}$-lift ${g_{d}:\mathbb{R}\to G}$ of $q_{d}$ with bounded body velocity. Let ${P}$ be a ${0_{Q}}$-navigation function and $\nu$ be a Riemannian metric on $Q$. For each (fixed) state ${\dot{q}\in TQ}$, let ${e^{q}:r\mapsto\Phi_{g_{d}(r)}^{-1}\big{(}q\big{)}}$ be a curve in $Q$ and ${X^{\dot{q}}:}$ ${r\mapsto\operatorname{d}\Phi_{g_{d}(r)}^{-1}(\dot{q})}$ be a vector field along $e^{q}$. Then, the control policy
$f_{q}(t,\dot{q})=-\operatorname{d}\Phi^{*}_{g_{d}^{-1}}\big{(}\operatorname{d}$ $P(e)+\nu^{\flat}(\dot{e})\,+$ (14)
$\kappa^{\flat}\circ\big{(}\nabla_{\dot{e}^{q}}(\dot{e}^{q}+2\,X^{\dot{q}})\big{)}(t)\big{)}$
achieves almost global asymptotic tracking of the reference and local exponential convergence of the tracking error.
Remark 4.
We use the “dummy variable” $r$ in $e\raisebox{3.5pt}{${\scriptstyle q}$}$ and $X\raisebox{3.5pt}{${\scriptstyle\dot{q}}$}$ to emphasize that $q$ and $\dot{q}$ are held fixed as $r$ varies (despite their dependence on $t$). This allows us to rigorously and intrinsically express (14) using only the standard formalism for covariant differentiation along curves. It follows from [7, Prop. 4.26] that (14) depends only on ${t\in\mathbb{R}}$ and ${\dot{q}\in TQ}$. Additionally, although the control policy (14) requires choosing a certain lift $g_{d}$, we will show that the qualitative closed-loop stability properties are ultimately independent of this choice.
Proof of Theorem 1.
The configuration error (7) is the “diagonal” curve of the smooth “family of curves” given by
$E:\mathbb{R}^{2}\to Q,\,(r,s)\mapsto\Phi\big{(}g_{d}(r)^{-1},q(s)\big{)},$ (15)
in the sense that ${e(t)=E(t,t)}$. We will use this observation to express the covariant derivative $\nabla_{\dot{e}}\dot{e}$ in terms of the respective contributions of the reference and actual trajectories.
Following [7, Ch. 6], the family of curves $E$ has transverse and main curves at each $s$ and $r$, given respectively by ${E^{s}:r\mapsto E(r,s)}$ and ${E_{r}:s\mapsto E(r,s).}$ Let $\mathfrak{X}(E)$ be the set of vector fields over $E$, i.e., maps ${(r,s)\mapsto V(r,s)}$ such that ${V(r,s)\in T_{E(r,s)}Q}$. For example, two such vector fields are the transverse and main velocity, ${\partial_{r}E:(r,s)\mapsto\dot{E}^{s}(r)}$ and ${\partial_{s}E:(r,s)\mapsto\dot{E}_{r}(s)}$. For any ${V\in\mathfrak{X}(E)}$, we may compute its “partial” covariant derivative along the transverse or main direction, operations denoted respectively by ${\operatorname{D}_{r},\,\operatorname{D}_{s}:\mathfrak{X}(E)\to\mathfrak{X}(E)}$. This operation is defined by restricting the vector field over the family of curves to a vector field along each transverse (resp. main) curve and computing the usual covariant derivative along that curve.
Covariant differentiation is a purely local operation [7, Lemma 4.1], and thus a local coordinate calculation can be used to show that because $e$ is the diagonal curve of $E$,
$\big{(}\nabla_{\dot{e}}\dot{e}\big{)}(t)$ $=\big{(}\operatorname{D}_{r}\partial_{r}\mathit{E}+2\operatorname{D}_{r}\partial_{s}\mathit{E}+\operatorname{D}_{s}\partial_{s}\mathit{E}\big{)}(t,t).$ (16)
where we have evaluated the right-hand side at $(r,s)=(t,t)$. We now aim to compute the terms on the right-hand side of (16). Observing that ${E^{s}(r)=e^{q(s)}}(r)$ and also that
$\partial_{s}E(r,s)=\operatorname{d}\Phi_{g_{d}(r)^{-1}}\big{(}\dot{q}(s)\big{)}=X^{\dot{q}(s)}(r),$ (17)
we may verify that
$\big{(}\operatorname{D}_{r}\partial_{r}E\big{)}(t,t)$ $=\big{(}\nabla_{\dot{E}^{t}}{\dot{E}^{t}}\big{)}(t)=\big{(}\nabla_{\dot{e}^{q}}{\dot{e}^{q}}\big{)}(t),$ (18)
$\big{(}\operatorname{D}_{r}\partial_{s}\mathit{E}\big{)}(t,t)$ $=(\nabla_{\dot{e}^{q}}X^{\dot{q}})(t).$ (19)
By [8, Thm. 5.70], the $\Phi$-invariance of $\kappa$ implies that $\nabla$ satisfies ${\nabla_{\operatorname{d}\Phi_{g}(X)}{\operatorname{d}\Phi_{g}(Y)}=\operatorname{d}\Phi_{g}\big{(}\nabla_{X}Y\big{)}}$ for all ${X,Y\in\mathfrak{X}(Q)}$ and ${g\in G}$. From this fact, (5), and (17), it follows that
$(\operatorname{D}_{s}\partial_{s}E)(t,t)=\big{(}\nabla_{\dot{E}_{t}}{\dot{E}_{t}}\big{)}(t)$ $=\operatorname{d}\Phi_{g_{d}(t)}^{-1}\circ\kappa^{\sharp}\big{(}f_{q}(t)\big{)},$ (20)
where ${t\mapsto f_{q}(t)}$ is the input force signal corresponding to ${t\mapsto q(t)}$. Substituting these results into (16), we obtain
$(\nabla_{\dot{e}}\dot{e})(t)=\nabla_{\dot{e}^{q}}(\dot{e}^{q}+2\,X^{\dot{q}})(t)+\operatorname{d}\Phi_{g_{d}(t)}^{-1}\circ\kappa^{\sharp}\big{(}f_{q}(t)\big{)}.$ (21)
Substituting (14) into (21) and using the $\Phi$-invariance of $\kappa$ to verify that ${(\operatorname{d}\Phi_{g}\circ\kappa^{\sharp})^{-1}=\operatorname{d}\Phi_{g}^{*}\circ\kappa^{\flat}}$ for any ${g\in G}$ ultimately yields the autonomous tracking error dynamics
$\nabla_{\dot{e}}\dot{e}=-\kappa^{\sharp}\big{(}\operatorname{d}P(e)+\nu^{\flat}(\dot{e})\big{)},$ (22)
which is a mechanical system with strict dissipation and a navigation function potential. Thus, it follows from [3, Thm. 2] that there exists an open dense set ${S\subseteq TQ}$ of full measure, such that for any ${t_{0}\in\mathbb{R}}$ (since (22) is autonomous), if ${\dot{e}(t_{0})\in S}$, then ${\dot{e}(t)\to 0_{TQ}}$ as ${t\to\infty}$, and local exponential stability $0_{TQ}$ for (22) follows from [8, Thm. 6.45].
Finally, it follows from [15, II.A.2] that ${\operatorname{dist}_{TQ}(u_{p},v_{r})\leq}$ ${\operatorname{dist}_{Q}(p,r)+\big{|}\big{|}\tau_{\gamma}(u_{p})-{v}_{r}\big{|}\big{|}_{\kappa}}$ for sufficiently close points ${p,r\in Q}$, where $\tau_{\gamma}$ denotes parallel transport along the shortest geodesic from $p$ to $r$. The $\Phi$-invariance of $\kappa$, the boundedness of the reference body velocity, and (9a)-(9b) can then be used with this inequality to verify that ${\dot{e}(t)\to 0_{TQ}}$ implies ${\operatorname{dist}_{TQ}\big{(}\dot{q}(t),\dot{q}_{d}(t)\big{)}\to 0}$. Since $\operatorname{d}\Phi_{g}$ is diffeomorphism, (9a) implies (for each ${t_{0}\in\mathbb{R}}$) the existence an open dense set ${S_{t_{0}}\subseteq TQ}$ of full measure such that if ${\dot{q}(t_{0})\in S_{t_{0}}}$, then ${\operatorname{dist}_{TQ}\big{(}\dot{q}(t),\dot{q}_{d}(t)\big{)}\to 0}$ as ${t\to\infty}$, proving the claim. ∎
V Specialization to $n$-Spheres and Lie Groups
We now specialize the controller proposed in Theorem 1 to the $n$-sphere and Lie group settings, in each case obtaining a concise and explicit expression for the control policy (14).
Corollary 1 (Almost Global Asymptotic Tracking on $\mathbb{S}^{n}$).
For any ${n\in\mathbb{N}}$, consider a fully-actuated mechanical system on ${(\mathbb{S}^{n},\Psi,\rho})$, a reference trajectory ${q_{d}:\mathbb{R}\to\mathbb{S}^{n}}$, and any ${0_{\mathbb{S}^{n}}}$-lift ${R_{d}:\mathbb{R}\to SO(n+1)}$ of $q_{d}$ with bounded body velocity. Then for any ${k_{P},k_{D}>0}$, the control policy
$f_{q}(t,\dot{q})=-k_{P}$ $\,{q_{d}}^{\textrm{T}}\big{(}q\,q^{\textrm{T}}-\mathds{1}\big{)}-k_{D}\big{(}\dot{q}^{\textrm{T}}+q^{\textrm{T}}{\dot{R}_{d}}\,{R_{d}}^{\textrm{T}}\big{)}$ (23)
$+\big{(}q^{\textrm{T}}{\ddot{R}_{d}}+2\,{\dot{q}}^{\textrm{T}}\dot{R}_{d}\big{)}\,{R_{d}}^{\textrm{T}}\big{(}q\,q^{\textrm{T}}-\mathds{1}\big{)}$
achieves almost global asymptotic tracking of the reference and local exponential convergence of the tracking error.
Proof.
It will suffice to apply Theorem 1 with the configuration error (12), the navigation function (3), and the dissipation metric $\nu=k_{D}\,\rho$. To show this, we compute
$\operatorname{d}P(e)$ $=(-k_{P}{0_{\mathbb{S}^{n}}})^{\textrm{T}}(\mathds{1}-e\,e^{\textrm{T}}\hskip 1.0pt),$ (24)
$\nu^{\flat}(\dot{e})$ $=k_{D}(q^{\textrm{T}}\,\dot{R}_{d}+\dot{q}^{\textrm{T}}\,R_{d}).$ (25)
Moreover, the covariant derivative along $\gamma$ of $X\in\mathfrak{X}(\gamma)$ is
$\nabla_{\dot{\gamma}}X:r\mapsto\big{(}\mathds{1}-\gamma(r)\,\gamma(r)^{\textrm{T}}\big{)}\,\dot{X}(r),$ (26)
and we have ${e^{q}(r)=R_{d}(r)^{\textrm{T}}\,q}$ and ${X^{\dot{q}}(r)=R_{d}(r)^{\textrm{T}}\,\dot{q}}$. Thus,
$\nabla_{\dot{e}^{q}}(\dot{e}^{q}+2\,X^{\dot{q}})(t)=\big{(}\mathds{1}-e\,e^{\textrm{T}}\hskip 1.0pt\big{)}\big{(}\ddot{R}_{d}{}^{\textrm{T}}\,q+2\,\dot{R}_{d}{}^{\textrm{T}}\,\dot{q}\big{)}.$ (27)
Finally, noting that ${\operatorname{d}\Psi^{*}_{R}(\omega)=\rho^{\flat}\circ\operatorname{d}\Psi_{R}^{-1}\circ\rho^{\sharp}(\omega)=\omega\,{R}^{\textrm{T}}}$ due to the $\Psi$-invariance of $\rho$, and also verifying that ${\big{(}\mathds{1}-e\,e^{\textrm{T}}\hskip 1.0pt\big{)}={R_{d}}^{\textrm{T}}\big{(}\mathds{1}-q\,q^{\textrm{T}}\big{)}{R_{d}}},$ we may compute (14) using all the previous calculations, simplifying to yield exactly (23). Thus, the claim follows immediately by Theorem 1. ∎
For any function ${P:G\to\mathbb{R}}$ on a Lie group $G$, we define a map ${\zeta_{P}:G\to\mathfrak{g}^{*}},\ {g\mapsto\operatorname{d}L_{g}^{*}\circ\operatorname{d}P(g)}$. Also, we denote the body velocities of trajectories $g$, $g_{d}$, and $e$ by by $\xi$, $\xi_{d}$, and $\xi_{e}$ respectively. The control policy we now recover uses the same configuration error as in [8, Thm. 11.29], although our different feedforward terms lead to almost global tracking. In that sense, our result is qualitatively more similar to [4], but they use a different configuration error (i.e., $e=g_{d}\,g^{-1}$).
Corollary 2 (Almost Global Asymptotic Tracking on a Lie Group).
For any Lie group $G$, consider a fully-actuated mechanical system on ${(G,L,\kappa_{\mathbb{I}})}$ and a reference trajectory ${g_{d}:\mathbb{R}\to G}$ with bounded body velocity. Let ${P}$ be a ${\mathds{1}}$-navigation function on $G$ and $\mathbb{D}$ be an inner product on $\mathfrak{g}$. Then for the virtual control ${\tau=\operatorname{d}L_{g}^{*}(f_{g})}$, the control policy
$\tau(t,g,\xi)$ $=-\zeta_{P}(e)-\mathbb{D}^{\flat}\xi_{e}\ +$ (28)
$\quad\ \,\mathbb{I}^{\flat}(\operatorname{Ad}_{e}^{-1}\dot{\xi}_{d}+[\xi,\xi_{e}])+\operatorname{ad}^{*}_{\xi_{e}}\mathbb{I}^{\flat}\xi_{e}-\operatorname{ad}^{*}_{\xi}\mathbb{I}^{\flat}\xi$
achieves almost global asymptotic tracking of the reference and local exponential convergence of the tracking error.
Proof.
It will suffice to apply Theorem 1 with the configuration error (13), the given navigation function, and the metric $\nu$ induced by $\mathbb{D}$ as in (2). Since $e^{g}(r)=L_{g_{d}(r)}^{-1}(g)$, we have
$\dot{e}^{g}(r)$ $=\operatorname{d}L_{e^{g}(r)}\big{(}-\operatorname{Ad}_{e^{g}(r)}^{-1}\big{(}\xi_{d}(r)\big{)}\big{)},$ (29)
$X^{\dot{g}}(r)=\operatorname{d}L_{g_{d}(r)}^{-1}(\dot{g})$ $=\operatorname{d}L_{e^{g}(r)}\big{(}\xi\big{)},$ (30)
so that altogether, we have
$(\dot{e}^{g}+2X^{\dot{g}})(r)$ $=\operatorname{d}L_{e^{g}(r)}\big{(}2\,\xi-\operatorname{Ad}_{e^{g}(r)}^{-1}\big{(}\xi_{d}(r)\big{)}\big{)}.$ (31)
The result [8, Thm. 5.40] implies that that for a curve $\gamma$ in $G$, a vector field ${X\in\mathfrak{X}(\gamma)}$, and curves $\upsilon,\eta$ in $\mathfrak{g}$ for which ${\dot{\gamma}(r)=\operatorname{d}L_{\gamma(r)}\big{(}\upsilon(r)\big{)}}$ and ${X(r)=\operatorname{d}L_{\gamma(r)}\big{(}\eta(r)\big{)}}$, we have
$\big{(}\nabla_{\dot{\gamma}}X\big{)}(r)=\operatorname{d}L_{\gamma(r)}\big{(}\dot{\eta}(r)+\nabla_{\mathbb{I}}\big{(}{\upsilon(r)},\eta(r)\big{)}\big{)},$ (32)
where ${\nabla_{\mathbb{I}}:\mathfrak{g}\times\mathfrak{g}\to\mathfrak{g}}$ is the bilinear map given by
$(\upsilon,\eta)\mapsto\tfrac{1}{2}[\upsilon,\eta]-\tfrac{1}{2}\mathbb{I}^{\sharp}\big{(}\operatorname{ad}_{\upsilon}^{*}\mathbb{I}^{\flat}\,\eta+\operatorname{ad}_{\eta}^{*}\mathbb{I}^{\flat}\,\upsilon\big{)}.$ (33)
Thus, after verifying that $\xi_{e}=\xi-\operatorname{Ad}_{e}^{-1}\xi_{d}$ and computing
$\tfrac{\operatorname{d}}{\operatorname{d}r}\big{(}2\,\xi$ $-\operatorname{Ad}_{e^{g}(r)}^{-1}\xi_{d}(r)\big{)}=-\operatorname{Ad}_{e^{g}(r)}^{-1}\dot{\xi}_{d}(r),$ (34)
we use (32) and the bilinearity of (33) to conclude that
$\nabla_{\dot{e}^{g}}(\dot{e}^{g}+2X^{\dot{g}})(t)$ (35)
$=-\operatorname{d}L_{e}\big{(}\operatorname{Ad}_{e}^{-1}\dot{\xi}_{d}+[\xi,\xi_{e}]+\mathbb{I}^{\sharp}(\operatorname{ad}_{\xi_{e}}^{*}\mathbb{I}^{\flat}\xi_{e}-\operatorname{ad}_{\xi}^{*}\mathbb{I}^{\flat}\xi)\big{)}.$
Finally, we define $f_{g}$ as in (14) and use (LABEL:evaluate_lie_group_covariant_term) to compute ${\tau=\operatorname{d}L_{g}^{*}(f_{g})}$, simplifying using the identities ${\operatorname{d}L_{g}^{*}\circ\operatorname{d}L_{g_{d}{}^{-1}}^{*}=\operatorname{d}L_{e}^{*}}$ and ${\operatorname{d}L_{e}^{*}\circ\nu^{\flat}=\nu^{\flat}\circ\operatorname{d}L_{e}^{-1}}$ to yield (28). Thus, the claim follows immediately by Theorem 1. ∎
VI Applications of the Method
Example 5 (The Axisymmetric Satellite).
Consider a free-floating satellite, modeled as an (underactuated) mechanical system on $(SO(3),L,\kappa_{\mathbb{J}})$ consisting of a rigid body with inertia tensor ${\mathbb{J}=\mathrm{diag}(J_{1},J_{2},J_{3})}$ with ${J_{1}=J_{2}}$ and control torques lying in a two-dimensional left-invariant codistribution corresponding to ${\operatorname{span}\{\hat{e}_{1},\hat{e}_{2}\}\subset\mathfrak{so}(3)^{*}}$, where ^${\hat{\cdot}:\mathbb{R}^{3}\to\mathfrak{so}(3)^{*}\cong\mathfrak{so}(3)}$ is the usual isomorphism satisfying $\hat{a}b=a\times b$. The satellite is equipped with a camera or antenna aligned with the $e_{3}$ axis, whose bearing is thus described by the output ${y:\dot{R}\mapsto R\,e_{3}}$. This system is invariant under the (left) action of $SO(2)$ on $SO(3)$ corresponding to a body-fixed rotation around $e_{3}$. By Noether’s theorem, the evolution around the symmetry axis is governed by conservation of momentum (i.e. ${\frac{\operatorname{d}}{\operatorname{d}t}\big{(}J_{3}\,{\Omega}_{3}(t)\big{)}=0}$). When ${\Omega_{3}(t_{0})=0}$, reduction by this symmetry (see [8, Thm. 5.83]) will yield a fully-actuated mechanical system on ${(\mathbb{S}^{2}=SO(3)/SO(2),\Psi,J_{1}\,\rho)}$. It is clear that output tracking for the original (underactuated) system on $SO(3)$ amounts to asymptotically tracking a state trajectory for the reduced (fully-actuated) system on $\mathbb{S}^{2}$. We apply Corollary 1 to synthesize the tracking controller, using the reductive decomposition ${\mathfrak{so}(3)=\mathfrak{f}\oplus\mathfrak{q}}$ with ${\mathfrak{f}=\operatorname{span}\{\hat{e}_{3}\}\cong\mathfrak{so}(2)}$ and ${\mathfrak{q}=\operatorname{span}\{\hat{e}_{1},\hat{e}_{2}\}}$ to compute a horizontally lifted reference in the manner of Proposition 2 (in fact, the principal connection is the “mechanical connection” [14, Sec. 3.10]), using local trivializations (see Remark 3) obtained from [16, Eqn. (29)-(30)]. Fig. 1 shows the trajectory tracking performance.
Example 6 (The Omnidirectional Aerial Robot [17]).
Consider an aerial robot consisting of a single rigid body and actuators capable of applying arbitrary wrenches. Compensating for gravity yields a fully-actuated mechanical system on $(\mathbb{R}^{3}\times SO(3),L,\kappa_{\mathbb{I}})$, where ${\mathbb{I}=}$ ${\mathrm{diag}(m\mathds{1}_{3\times 3},\mathbb{J}_{3\times 3})}$. We use Corollary 2 with (4) to track trajectories (see Fig. 2).
Figure 2: Almost global asymptotic trajectory tracking on ${\mathbb{R}^{3}\times SO(3)}$.
VII Discussion
Of the $n$-spheres, only $\mathbb{S}^{1}$ and $\mathbb{S}^{3}$ admit a Lie group structure, and thus methods like [4] are not applicable to tracking on, e.g., $\mathbb{S}^{2}$. As pointed out in [1], methods which do not rely on reduction to regulation often fail to achieve or certify almost global convergence, in large part because they cannot benefit from [3, Thm. 2]. Other smooth tracking controllers on $\mathbb{S}^{2}$ (e.g. [18] or [19, Sec. III]) do not guarantee convergence from almost every initial state in $TQ$, rather ensuring convergence from a smaller region that, e.g., may contain the zero tangent vector over almost every point in $Q$ (not $TQ$), sometimes leading to an abuse of the term “almost global” in regards to asymptotic stability. We also note that the differential properties of a state-valued tracking error are important, since it was the surjectivity of the partial derivative of the tracking error appearing in (21) that enabled the feedforward cancellation of other terms and the injection of arbitrary dissipation and potential, regardless of the system considered (a difficulty of the approach proposed in [1]).
Example 5 demonstrates applicability to output tracking for certain underactuated systems. Moreover, hierarchical controllers for underactuated systems (e.g., [20]) often rely on the control of subsystems that “look” fully-actuated, and the identification of a “geometric flat output” [16] can often yield such a decomposition wherein the subsystems evolve on homogeneous spaces that may not be Lie groups. Also, the almost global asymptotic stability of a hierarchical controller with subsystems of the form proposed in Theorem 1 can be certified compositionally using only the properties of the subsystems and the “interconnection term” [21], thanks to the dissipative mechanical form of the error dynamics (22). For these reasons, we believe the proposed method has substantial implications in the systematic synthesis of controllers for a broad class of underactuated systems.
VIII Conclusion
We propose a systematic, unified tracking controller for fully-actuated mechanical systems evolving on homogeneous Riemannian manifolds, with exactly zero probability of nonconvergence from a randomly selected initial state. We apply the method to systems with two different configuration manifolds. Conceptually, our results illustrate that it is the transitivity of a Lie group’s action on the configuration manifold (and not the absence of fixed points) that is useful for tracking control via error regulation in mechanical systems.